# Online estimation of the inverse of the Hessian for stochastic optimization with application to universal stochastic Newton algorithms

### v1

@@ 5.2 Comparison with Riccati Newton
The objective of this section is to demonstrate the comparable performance of USNA and UWASNA when contrasted with SNA and WASNA, particularly in scenarios where the use of the Riccati formula is applicable to recursively compute the inverse of an estimator of the Hessian. Let us begin by revisiting this context. If we can estimate the Hessian matrix $H=\nabla^{2}G(\theta)$ by an estimator of the form $S_{n}/n$ with $S_{n}$ defined by $S_{n}=\sum_{j=1}^{n}\varphi_{j}\varphi_{j}^{T}$ where $(\varphi_{n})_{n\geq 1}$ is a sequence of random vectors of ${\mathbb{R}}^{d}$, then, thanks to the Riccati formula, we can recursively calculated matrix $S_{n}^{-1}$ for any $n\geq 1$:
@@ 5.2.2 Geometric Median
In this context, algorithm (11)-(13) can be used to estimate $m$ taking $\nabla_{h}g(X_{n},m_{n-1}^{SN})=-\frac{X_{n}-m_{n-1}}{\left\|X_{n}-m_{n-1}\right\|}$ and $\varphi_{n}=\nabla_{h}g(X_{n},m_{n-1}^{SN}+\alpha_{n}Z_{n})-\nabla_{h}(X_{n},m_{n-1}^{SN})$, where $\alpha_{n}=\frac{1}{n\ln{(n+1)}}$ and $(Z_{n})_{n\geq 1}$ is a sequence of independent standard Gaussian vectors.
@@ 5.3.1 Spherical Distribution
In this paragraph, we focus on the estimation of the parameters of a spherical distribution (Godichon-Baggioni and Portier,, 2017). The aim of the task is to fit a sphere onto a 3D point cloud with noise. In this context, we assume that the observations represent independent realizations of a random vector $X$, which is defined as
@@
Figure 3 illustrates the comparison between the performances of USNA, UWASNA and ASGD in terms of the mean squared error (MSE) of the estimated parameters. Throughout the simulations, UWASNA amd USNA demonstrate superior performance in accurately estimating the parameters of the spherical distribution when compared to the gradient-based method. Additionally, the Hessian matrix $H$ of the model can be explicitly calculated (Godichon-Baggioni and Portier,, 2017), one has
$H=\begin{pmatrix}I_{3}-\frac{2}{3}I_{3}\mathbb{E}\left[W\right]\left[W^{-1}\right]&0\\ 0&1\end{pmatrix}.$
@@
Now we focus on the estimation of p-means of a multivariate distribution (Fréchet,, 1948). We consider a random vector $X$ of $\mathbb{R}^{d}$ with $d=40$, where $X\sim\mathcal{N}\left(0,\bf{\Sigma}\right)$ with $\bf{\Sigma}_{i,j}=|i-j|^{0.5}$. The p-mean $m$ of $X$ is defined as the minimizer of the functional $G_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ given for all $h\in\mathbb{R}^{d}$ by:
@@ 5.4 Application to real data
We apply the algorithms to ”COVTYPE” dataset, a well-known dataset used for classification tasks (Blackard and Dean,, 1999; Schmidt et al.,, 2017; Toulis et al.,, 2016). The original study was based on a dataset comprising 581,011 observations and 54 predictors. The primary objective was to predict the cover type of forests within Roosevelt National Park. In our current investigation, we will narrow our focus to the most prevalent category of the target variable, which is ”Spruce/Fir,” accounting for 49% of the observations. As a result, we convert the ”covertype” variable into a binary form, with the ”fir” category assigned a value of 1, while all other categories are assigned a value of 0. The objective is to use logistic regression to predict the variable ”covertype”. We split the dataset into training (80%) and test (20%) sets.
We implement the UWASNA, USNA, WASNA and SNA on the training set. As a baseline, we also apply a first-order algorithm ASGD on it. We evaluate the performance of each algorithm by calculating the accuracy of each one on both training set and test set. The results are summarized in Table 1.
UWASNA USNA WASNA SNA ASGD
Training Accuracy(%) 75.63 75.31 75.52 75.38 74.54
Test Accuracy(%) 75.61 75.34 75.50 75.33 74.64
Table 1: Accuracy of UWASNA, USNA, WASNA, SNA and ASGD algorithms on ”COVTYPE” dataset.
We observe that UWASNA, USNA, WASNA and SNA demonstrate similar performances, and they achieve higher accuracies on both training set and test set compared to ASGD. By successfully applying USNA and UWASNA on this dataset, we illustrate their practicality in real-world applications.
@@ Conclusion
In this study, we thoroughly examined the stochastic optimization problem, primarily focusing on accurately estimating the unknown parameter. Our significant contribution is the introduction of a direct method to estimate the inverse of the Hessian matrix. Instead of the traditional approach, which estimates the Hessian matrix first, we directly addressed its inverse, using the Robbins-Monro procedure.
This approach led us to develop the Universal Weighted Averaged Stochastic Newton Algorithm. Through our extensive testing, we found that our newly proposed methods are both efficient and robust. When compared with standard first and second-order algorithms, our method consistently performed well. In certain scenarios, it even outperformed these conventional algorithms.
In summary, our findings emphasize the potential of second-order methods in optimization. Our approach to directly estimating the Hessian matrix’s inverse represents a significant advancement, combining simplicity with efficiency.
@@ 6 Proofs
For the sake of simplicity, in the following we denote $T_{n}:=\nabla^{2}_{h}g(X_{n},\hat{\theta}_{n-1})Z_{n}Z_{n}^{T}$. Recall that $\left\|Z_{n}\right\|$ is bounded, i.e there is $M$ such that $\left\|Z\right\|\leq M$. In addition, let us recall that if $\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}$, it then follows that $\left\|T_{n+1}\right\|_{op}\leq\beta_{n+1}$.
@@ 6.1 Study on the largest eigenvalue of $A_{n}$
[…]
@@
Define $W_{n}:=A_{n-1}T_{n}^{T}+T_{n}A_{n-1}$. By definition of $A_{n}$,
$\left\|A_{n+1}\right\|_{F}^{2}$ $=\left\|A_{n}\right\|_{F}^{2}-2\gamma_{n+1}\braket{A_{n},W_{n+1}-2I_{d}}_{F}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}$
@@
Therefore,
$\mathbb{E}\left[\left\|A_{n+1}\right\|_{F}^{2}|\mathcal{F}_{n}\right]$ $=\left\|A_{n}\right\|_{F}^{2}-2\gamma_{n+1}\braket{A_{n},\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}}_{F}$
@@
Assumption (A5) ensures that
$\mathbb{E}$ $\left[\left\|W_{n+1}-2I_{d}\right\|^{2}_{F}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}|\mathcal{F}_{n}\right]$
$\leq 8\left(\mathbb{E}\left[\left\|\nabla^{2}_{h}g(X_{n+1},\hat{\theta}_{n})\right\|_{F}^{2}\right]\mathbb{E}\left[\left\|Z_{n+1}\right\|^{4}\right]\left\|A_{n}\right\|^{2}_{F}+d\right)$
$\leq 8\left(C_{q}^{2/q}\left\|A_{n}\right\|^{2}_{F}+d\right).$
Since $\left\|Q_{n+1}\right\|$ = $\left\|Q_{n+1}\right\|^{q}\left\|Q_{n+1}\right\|^{1-q}$ with $q>1$, for any $\zeta_{n}>0$,
@@
$+\frac{\gamma_{n+1}}{\zeta_{n+1}}\left\|A_{n}\right\|^{2}_{F}+\zeta_{n+1}d$
@@
Since $\nabla^{2}G(\hat{\theta}_{n})$ and $A_{n}$ are positive, one has for any $\zeta_{n}>0$,
$-\gamma_{n+1}\braket{A_{n},\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}}_{F}$ $\leq\gamma_{n+1}\braket{A_{n},2I_{d}}$
@@
Setting $\zeta_{n}=n^{1-\gamma}\ln n^{1+\delta}$ with $\delta>0$, we can apply Lemma 6.2 with $V_{n}=\left\|A_{n}\right\|_{F}^{2}$, and $a_{n}=\zeta_{n}^{2}$. One then obtains $\left\|A_{n}\right\|_{F}^{2}=o\left(\zeta_{n}^{2}\right)=o\left(n^{2-2\gamma}\ln n^{2+2\delta}\right)$, so that $\lambda_{\max}(A_{n})=o\left(n^{1-\gamma}\ln n^{1+\delta}\right)a.s.$
@@
The aim is to provide an initial rate of convergence for $A_{n}$. A more refined or faster rate will be established later. First, note that
$A_{n+1}-H^{-1}$ $=A_{n}-H^{-1}-\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}$
$=A_{n}-H^{-1}-\gamma_{n+1}(W_{n+1}-2I_{d})+\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}$
$=A_{n}-H^{-1}-\gamma_{n+1}\left(\mathbb{E}\left[W_{n+1}|\mathcal{F}_{n}\right]-2I_{d}\right)+\gamma_{n+1}\xi_{n+1}$

### v3

@@ 5.2 Comparison with Riccati Newton
The objective of this section is to demonstrate the comparable performance of USNA and UWASNA when contrasted with SNA and WASNA, particularly in scenarios where the use of the Riccati formula is applicable to recursively compute the inverse of an estimator of the Hessian. Let us begin by revisiting this context. If we can estimate the Hessian matrix $H=\nabla^{2}G(\theta)$ by an estimator of the form $S_{n}/n$ with $S_{n}$ defined by $S_{n}=\sum_{j=1}^{n}\varphi_{j}\varphi_{j}^{T}$ where $(\varphi_{n})_{n\geq 1}$ is a sequence of random vectors of ${\mathbb{R}}^{d}$, then, thanks to the Riccati formula   (also called Sherman-Morrison formula), we can recursively calculate matrix $S_{n}^{-1}$ for any $n\geq 1$:
@@ 5.2.2 Geometric Median
In this context, algorithm (11)-(13) can be used to estimate $m$ taking $\nabla_{h}g(X_{n},m_{n-1}^{SN})=-\frac{X_{n}-m_{n-1}}{\left\|X_{n}-m_{n-1}\right\|}$ and $\varphi_{n}=\frac{\sqrt{\left\|X_{n}-m_{n-1}\right\|}}{\alpha_{n}}\left(\nabla_{h}g(X_{n},m_{n-1}^{SN}+\alpha_{n}Z_{n})-\nabla_{h}(X_{n},m_{n-1}^{SN})\right)$, where $\alpha_{n}=\frac{1}{n\ln{(n+1)}}$ and $(Z_{n})_{n\geq 1}$ is a sequence of independent standard Gaussian vectors.
@@ 5.3.1 Spherical Distribution
In this paragraph, we focus on the estimation of the parameters of a spherical distribution (Godichon-Baggioni and Portier,, 2017). This is an interesting example since we are able to explicitely calculate $\Sigma^{-1/2}$. Nevertheless we are not able to verify that our assumptions are satisfied in this context. The aim of the task is to fit a sphere onto a 3D point cloud with noise. In this context, we assume that the observations represent independent realizations of a random vector $X$, which is defined as
@@
Figure 3 illustrates the comparison between the performances of USNA, UWASNA and ASGD in terms of the mean squared error (MSE) of the estimated parameters. Throughout the simulations, UWASNA and USNA demonstrate superior performance in accurately estimating the parameters of the spherical distribution when compared to the gradient-based method. Additionally, the Hessian matrix $H$ of the model can be explicitly calculated (Godichon-Baggioni and Portier,, 2017), one has
$H=\begin{pmatrix}I_{3}-\frac{2}{3}I_{3}\mathbb{E}\left[W\right]\mathbb{E}\left[W^{-1}\right]&0\\ 0&1\end{pmatrix}.$
@@
Now we focus on the estimation of p-means of a multivariate distribution (Fréchet,, 1948). We consider a random vector $X$ of $\mathbb{R}^{d}$ with $d=40$, where $X\sim\mathcal{N}\left(0,\mathbf{\Sigma}\right)$ with $\mathbf{\Sigma}_{i,j}=|i-j|^{0.5}$. The p-mean $m$ of $X$ is defined as the minimizer of the functional $G_{p}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ given for all $h\in\mathbb{R}^{d}$ by:
@@ 5.4 Application to real data
We now compare the performance of various algorithms across several real-world datasets. Table 1 summarizes the main characteristics of the different dataset used in our experiments.
Dataset # Features Training Set Size Testing Set Size
Epsilon 2000 400,000 100,000
Mushrooms 112 6,499 1,625
Web Page 300 3,470 46,279
MNIST 784 11,378 1,924
Phishing 68 5,527 5,528
Adult 123 32,561 16,281
COVTYPE 54 290,506 290,505
Table 1: Key characteristics of the datasets used in this study.
Epsilon (Chang and Lin,, 2011) is a large-scale binary classification dataset from the LIBSVM collection. Mushrooms (Lichman et al.,, 2013) consists of morphological attributes of mushrooms used to determine their edibility. Web Page (Platt,, 1999) involves webpage classification and is widely used in SVM-based text classification research. MNIST (LeCun et al.,, 1998) contains images of handwritten digits (0–9); for binary classification, we use only the digits 2–5. Phishing (Mohammad and McCluskey,, 2012) is designed to detect malicious websites and consists entirely of categorical features, which we encode in binary form. Adult (Dua et al.,, 2017) contains census data for income prediction and includes multiple categorical variables converted into binary indicators. COVTYPE (Blackard,, 1998) originally includes multiple forest cover types; in this study, we focus on distinguishing “Spruce/Fir” (labeled as 1) from all other categories (labeled as 0). These datasets are frequently adopted to benchmark binary classifiers (Toulis and Airoldi,, 2017; Juan et al.,, 2016; Yuan et al.,, 2012). While Epsilon, Web Page, MNIST, and Adult already include predefined training and testing sets, we partition COVTYPE, Mushrooms, and Phishing equally to create our own splits. Our main objective is to apply a logistic regression model to predict the binary response variable for each of the seven datasets.
We implement the UWASNA, USNA, WASNA and SNA on the training set. As a baseline, we also apply first-order algorithms SGD and AdaGrad on it. Since these methods are designed to be online (recursive), we process the training set in a single pass and do not use any stopping criterion, unlike iterative (offline) methods. We evaluate the performance of each algorithm by calculating the accuracy (percentage of good classifications) of each one on both training set and test set. The results are summarized in Table 2.
Dataset Set SNA WASNA USNA UWASNA AdaGrad SGD
Epsilon Train 89.94 90.05 84.51 79.83 87.18 71.79
Test 89.68 89.81 84.43 79.61 87.05 71.66
Mushrooms Train 99.98 100.00 99.24 98.50 99.21 97.46
Test 99.83 100.00 98.87 98.84 98.82 97.02
Web Page Train 98.56 98.04 96.92 97.00 97.00 96.92
Test 98.02 97.58 97.04 97.05 97.04 97.04
MNIST Train 94.28 93.08 96.36 96.53 97.21 96.99
Test 94.59 92.88 96.99 96.36 97.35 97.04
Phishing Train 93.56 93.87 92.80 92.62 92.82 85.67
Test 93.38 93.89 92.58 92.42 92.84 85.60
Adult Train 84.92 84.73 84.80 84.71 84.76 83.98
Test 85.03 85.00 84.94 85.17 85.09 84.10
COVTYPE Train 75.73 75.70 75.38 75.63 75.56 75.52
Test 75.86 75.78 75.50 75.69 75.67 75.65
Table 2: Train and Test accuracy (in %) for SNA, WASNA, USNA, UWASNA, AdaGrad, and SGD on various datasets.
We can observe that UWASNA, USNA, WASNA and SNA demonstrate similar performances, and they achieve higher accuracies on both training set and test set compared to SGD. By successfully applying USNA and UWASNA on this dataset, we illustrate their practicality in real-world applications.
@@ Conclusion
In this article, we introduce a novel second-order stochastic algorithm, referred to as universal Newton, and designed to estimate the unique minimizer of a convex function. Unlike the classical approach, which involves estimating and explicitly inverting the Hessian matrix, our method directly and recursively estimates its inverse using a Robbins-Monro algorithm. By incorporating this estimation into the optimization process, we develop the universal stochastic Newton algorithm, along with its averaged and weighted variants. We establish the convergence of these algorithms and their asymptotic efficiency. Numerical experiments on simulated data highlight the advantages of our approach over classical methods, while tests on real-world datasets confirm its applicability. Compared to standard first- and second-order algorithms, our method consistently yields comparable results and, in some scenarios, even outperforms them.
Acknowledgments
The authors would like to thank Guillaume Sallé for his help in improving the proofs.
@@ 6 Proofs
6.1 Notations and preliminary definitions
[…]
  • $(Z_{n})_{n\geq 1}$ is a sequence of independent and identically distributed random vectors of $\mathbb{R}^{d}$, independent of the sequence $(X_{n})_{n\geq 1}$, and such that for any $n\geq 1$, $\|Z_{n}\|\leq M$ and $\mathbb{E}\left[Z_{n}Z_{n}^{T}\right]=I_{d}$;
  • $\hat{\theta}_{n}$ is an estimate of unknown parameter $\theta$;
  • $T_{n+1}:=\nabla^{2}_{h}g(X_{n+1},\hat{\theta}_{n})Z_{n+1}Z_{n+1}^{T}$ ;
  • $A_{n}$, the estimate of $H^{-1}$, is recursively defined for all $n\geq 1$ by:
    $P_{n+1}$ $=A_{n}Z_{n}$
    $Q_{n+1}$ $=\nabla_{h}^{2}g(X_{n+1},\widehat{\theta}_{n})Z_{n+1}$
    $A_{n+1}$ $={\Pi_{\beta_{n+1}^{\prime}}}\left(A_{n}-\gamma_{n+1}\left(P_{n+1}Q_{n+1}^{T}+Q_{n+1}P_{n+1}^{T}-2\,I_{d}\right)\mathbf{1}_{\{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}\}}\right),$
    where $\Pi_{\beta_{n}^{\prime}}$ is the projection onto the ball of radius $\beta_{n+1}^{\prime}$ (with respect to the Frobenius norm);
  • $W_{n+1}:=A_{n}T_{n+1}^{T}+T_{n+1}A_{n}$.
Finally, in the proof of Theorems 4.1 and 4.2, we will use the following notations
$\xi_{n+1}$ $=-W_{n+1}+\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})$
$\Pi_{n+1}^{\perp}$ $=A_{n+1}-A_{n}+\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}$
$+\left(\nabla^{2}G(\hat{\theta}_{n})-H\right)\left(A_{n}-H^{-1}\right)+\left(A_{n}-H^{-1}\right)\left(\nabla^{2}G(\hat{\theta}_{n})-H\right)-\gamma_{n+1}^{2}H(A_{n}-H^{-1})H,$
$s^{\prime}_{n}$ $=\mathbb{E}\left[\left(T_{n+1}H^{-1}+H^{-1}T_{n+1}-2I_{d}\right)\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}|\mathcal{F}_{n}\right].$
6.2 Positivity of $A_{n}$
Let us prove by induction that $A_{n}$ is positive semi-definite for any $n\geq 0$. Let us denote by $\tilde{A}_{n+1}:=A_{n}-\gamma_{n+1}\left(A_{n}T_{n+1}^{T}+T_{n+1}A_{n}-2I_{d}\right)\mathbf{1}_{\lVert Q_{n+1}\rVert\leq\beta_{n}}$. If $\lVert Q_{n+1}\rVert>\beta_{n}$, then $\tilde{A}_{n+1}=A_{n}$ and is so positive (as well as the projection $A_{n+1}$). Conversely, if $\lVert Q_{n+1}\rVert\leq\beta_{n}$, we have
$\tilde{A}_{n+1}$ $=A_{n}-\gamma_{n+1}\left(A_{n}T_{n+1}^{T}+T_{n+1}A_{n}-2I_{d}\right)$
$=\left(I_{d}-\gamma_{n+1}T_{n+1}\right)A_{n}\left(I_{d}-\gamma_{n+1}T_{n+1}\right)^{T}-\gamma_{n+1}^{2}T_{n+1}A_{n}T_{n+1}^{T}+2\gamma_{n+1}I_{d}$
Then, since $\gamma_{n+1}^{2}\beta_{n+1}^{2}\beta_{n+1}^{\prime}\leq\gamma_{n+1}$,
$\lambda_{\min}\left(\tilde{A}_{n+1}\right)$ $\geq\lambda_{\min}\left(\left(I_{d}-\gamma_{n+1}T_{n+1}\right)A_{n}\left(I_{d}-\gamma_{n+1}T_{n+1}\right)^{T}\right)-\gamma_{n+1}^{2}\lambda_{\max}\left(T_{n+1}A_{n}T_{n+1}^{T}\right)+2\gamma_{n+1}$
$\geq 0-\gamma_{n+1}^{2}\beta_{n+1}^{2}\beta_{n+1}^{\prime}+2\gamma_{n+1}$
$\geq\gamma_{n+1}.$
Then, $\tilde{A}_{n+1}$ is positive (and its projected version $A_{n+1}$ too).
@@ 6.3 Study on the largest eigenvalue of $A_{n}$
$\lambda_{\max}(A_{n})=o\left(n^{1-\gamma}\ln n^{1+\delta}\right)a.s.\quad\text{and}\quad\lambda_{\max}(A_{n,\tau})=o\left(n^{1-\gamma}\ln n^{1+\delta}\right)a.s.$
@@
Define $W_{n}:=A_{n-1}T_{n}^{T}+T_{n}A_{n-1}$. By definition of $A_{n}$, and since the projection is $1$-Lipschitz,
$\left\|A_{n+1}\right\|_{F}^{2}$ ${\leq}\left\|A_{n}\right\|_{F}^{2}-2\gamma_{n+1}\braket{A_{n},W_{n+1}-2I_{d}}_{F}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}$
@@
Considering $\mathcal{F}_{n}=\sigma\left\{X_{1},...,X_{n},Z_{1},...,Z_{n}\right\}$, remark that
$\mathbb{E}\left[\braket{A_{n},W_{n+1}-2I_{d}}_{F}|\mathcal{F}_{n}\right]=\braket{A_{n},\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}}_{F}.$
Since $\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}=1-\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}$ it comes
$\mathbb{E}\left[\left\|A_{n+1}\right\|_{F}^{2}|\mathcal{F}_{n}\right]$ ${\leq}\left\|A_{n}\right\|_{F}^{2}-2\gamma_{n+1}\braket{A_{n},\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}}_{F}$
@@
Since $\|A+B\|_{F}^{2}\leq 2\|A\|_{F}^{2}+2\|B\|_{F}^{2}$ (used twice) and $\|AB\|_{F}\leq\|A\|_{F}\|B\|_{F}$, and since $\|Z_{n+1}\|\leq M$, Assumption (A4) ensures that
$\mathbb{E}\left[\left\|W_{n+1}-2I_{d}\right\|^{2}_{F}\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}|\mathcal{F}_{n}\right]$ $\leq 2\mathbb{E}\left[\left\|W_{n+1}\right\|_{F}^{2}|\mathcal{F}_{n}\right]+2\left\|2I_{d}\right\|_{F}^{2}$
$\leq 8\left(\mathbb{E}\left[\left\|\nabla^{2}_{h}g(X_{n+1},\hat{\theta}_{n})\right\|_{F}^{2}{\left\|Z_{n+1}\right\|^{4}}\left\|A_{n}\right\|^{2}_{F}|\mathcal{F}_{n}\right]+d\right)$
$\leq 8\left(C_{q}^{2/q}M^{4}\left\|A_{n}\right\|^{2}_{F}+d\right).$
Since $\left\|Q_{n+1}\right\|$ = $\left\|Q_{n+1}\right\|^{q}\left\|Q_{n+1}\right\|^{1-q}$ with $q>1$, using Cauchy-Schwarz inequality one has $\langle A_{n},W_{n+1}\rangle_{F}\leq\|A_{n}\|_{F}\|W_{n+1}\|_{F}$, and since for any $\zeta_{n+1}>0$, $\langle A_{n},I_{d}\rangle_{F}\leq\frac{1}{2\zeta_{n+1}}\|A_{n}\|_{F}^{2}+\frac{1}{2}\zeta_{n+1}\|I_{d}\|_{F}^{2}$,
@@
$+\frac{\gamma_{n+1}}{\zeta_{n+1}}\left\|A_{n}\right\|^{2}_{F}+\zeta_{n+1}\gamma_{n+1}d$
@@
Since $\nabla^{2}G(\hat{\theta}_{n})$ and $A_{n}$ are positive, one has for any $\zeta_{n+1}>0$,
$-\gamma_{n+1}\braket{A_{n},\nabla^{2}G(\hat{\theta}_{n})A_{n}+A_{n}\nabla^{2}G(\hat{\theta}_{n})-2I_{d}}_{F}$ $\leq\gamma_{n+1}\braket{A_{n},2I_{d}}_{F}$
@@
Setting $\zeta_{n}=n^{1-\gamma}\ln n^{1+\delta}$ with $\delta>0$, we can apply Lemma B.2 with $V_{n}=\left\|A_{n}\right\|_{F}^{2}$, and $a_{n}=\zeta_{n}^{2}$. One then obtains $\left\|A_{n}\right\|_{F}^{2}=o\left(\zeta_{n}^{2}\right)=o\left(n^{2-2\gamma}\ln n^{2+2\delta}\right)$, so that $\lambda_{\max}(A_{n})=o\left(n^{1-\gamma}\ln n^{1+\delta}\right)a.s.$
@@
6.4.1 Rate of convergence of $A_{n}$
The aim is to provide an initial rate of convergence for $A_{n}$. A more refined or faster rate will be established later. First, recalling that $\tilde{A}_{n+1}=A_{n}-\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}$, and denoting $\Pi_{n+1}^{\perp}=A_{n+1}-\tilde{A}_{n+1}$, observe that
$A_{n+1}-H^{-1}$ $=A_{n}-H^{-1}-\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|\leq\beta_{n+1}}{+\Pi_{n+1}^{\perp}}$
$=A_{n}-H^{-1}-\gamma_{n+1}(W_{n+1}-2I_{d})+\gamma_{n+1}(W_{n+1}-2I_{d})\mathbf{1}_{\left\|Q_{n+1}\right\|\left\|Z_{n+1}\right\|>\beta_{n+1}}{+\Pi_{n+1}^{\perp}}$
$=A_{n}-H^{-1}-\gamma_{n+1}\left(\mathbb{E}\left[W_{n+1}|\mathcal{F}_{n}\right]-2I_{d}\right)+\gamma_{n+1}\xi_{n+1}{+\Pi_{n+1}^{\perp}}$